A bond issued by a corporation with a provision that allows the issuer to repurchase the...

  1. A bond issued by a corporation with a provision that allows the issuer to repurchase the bond at a premium over par after a fixed time interval is called a:
  1. Treasury bond
  2. Debenture
  3. Development bond
  4. Callable bond
  1. In general, which of the following bonds carry the highest level of risk?
    1. Treasury Notes
    2. Corporate Bonds
    3. Municipal bonds
    4. Treasury bonds
  1. An 8% coupon bond maturing in 5 years has a yield to maturity of 10% and makes coupon payments semi-annually. What was the market rate when the bond was first issued?
    1. 8%
    2. 10%
    3. 12%
    4. 15%
  1. Vision Corp. offers a semiannual, 8% bond while Morning Corp. offers a similar, annual bond within the same risk class. The required rate on Vision is 4.5% and that of Morning is 5%. Which bond offers today the better yield?
    1. Vision Corp
    2. Morning Corp
    3. Both offer the same yield
    4. Neither offers a reasonable yield
Answer #1

1) Callable bond

2) Corporate bonds

3) The market rate at issue date will be the same as the coupon rate, which is 8%

4) Morning Corp offers a better yield today, Let's assume face value to be 1000 and years to maturity to be 10 years, the current price of each bond will be:

=PV(4.5%/2,10*2,8%/2*1000,1000) =1279.36 & =PV(5%,10,8%*1000,1000) =1231.65

Current yield will be =80/1279.36 =6.25% & =80/1231.65 = 6.50%

So Morning corp has better current yield.